OpenStudy (anonymous):

Solve the equation. 7 + 8 ln x = 12

OpenStudy (anonymous):

start with \[8\ln(x)=5\] then \[\ln(x)=\frac{5}{8}\] and finally \[x=e^{\frac{5}{8}}\]